OpenStudy (anonymous):

You are decorating a cylindrical hat box to give to your grandmother as a gift. You want to paste family photos onto the curved part of the box. The height of the box is 6 inches and the radius is 7 inches. If none of the photos overlap, what is the area of photographs you need? Use three point one four for pi. one hundred thirty-one point eight eight square inches two hundred sixty-three point seven six square inches seven hundred ninety-one point two eight square inches nine hundred twenty-three point one six sq

OpenStudy (anonymous):

@jordanloveangel @iGreen @cwrw238

OpenStudy (jordanloveangel):

1 sec

OpenStudy (jordanloveangel):

(I believe the word you're looking for is the Greek letter "pi", not the dessert frequently served with ice cream.) Area of a cylinder = πDh = 2πRh = 2(3.14)(6)(8) = 301.44 square inches.

OpenStudy (jordanloveangel):

r = 6'' h = 8'' w = 2*pi*r = 37.68'' a = h*w = 301.44 in^2

OpenStudy (jordanloveangel):

did it help ?

OpenStudy (anonymous):

its not an answer choice but i seen that number before abc or d just in case i just need glasses lol @jordanloveangel

OpenStudy (jordanloveangel):

kk hold on thn

OpenStudy (jordanloveangel):

jus find out the curved surface area : 2πr2h

OpenStudy (jordanloveangel):

they want you to cover this part, the side, of the cylinder the surface area of a right circular cylinder is A=2πrh where r is the radius and h is the height of the cylinder he formula above is only for the side I meant* which is what you want to cover with photos (the curved part)

OpenStudy (anonymous):

i already have the answer

OpenStudy (jordanloveangel):

kk does this help

OpenStudy (anonymous):

thank tho

OpenStudy (jordanloveangel):

yeah np